OpenStudy (leahtejeda):

Simplify 8 over the quantity of 2 plus 2i. −2 + 2i −2 − i −2 + i 2 − 2i

OpenStudy (campbell_st):

so is this to do with complex numbers... i^2 = -1?

OpenStudy (campbell_st):

if the problem is \[\frac{8}{2 + 2i}\] you can multiply by the conjugate pair so it becomes \[\frac{8}{2 + 2i} \times \frac{2 -2i}{2 - 2i}\] you should notice this is the difference of 2 squares in the denominator, so will rationalize the denominator hope it helps

OpenStudy (anonymous):

2-2i that is your answer